### v1

# Proximal Gradient Dynamics: Monotonicity, Exponential Convergence, and Applications

Abstract
In this letter, we study the proximal gradient dynamics. This recently-proposed continuous-time dynamics solves optimization problems whose cost functions are separable into a nonsmooth convex and a smooth component. First, we show that the cost function decreases monotonically along the trajectories of the proximal gradient dynamics. We then introduce a new condition that guarantees exponential convergence of the cost function to its optimal value, and show that this condition implies the proximal Polyak-Łojasiewicz condition. We also show that the proximal Polyak-Łojasiewicz condition guarantees exponential convergence of the cost function. Moreover, we extend these results to time-varying optimization problems, providing bounds for equilibrium tracking. Finally, we discuss applications of these findings, including the LASSO problem, quadratic optimization with polytopic constraints, and certain matrix based problems.
I Introduction
Optimization problems are central to various domains of engineering and data science. Many of these problems can be framed as minimizing a smooth cost function combined with a nonsmooth but convex regularizer that promotes structure in the solution. For example, the addition of the $\ell_{1}$ norm as a regularizer induces sparsity, a key principle in compressed sensing and sparse coding [3]. In other cases, convex constraints can be incorporated using indicator functions. A common approach for solving such problems, where the cost function is separable into a smooth, possibly nonconvex, and a nonsmooth convex component, is the proximal gradient algorithm [16]. This (discrete-time) algorithm gained prominence in sparse encoding applications [3]. A list of widely used regularizers and their associated proximal operators is available in [2, Chapter 10].
Analyzing the continuous-time versions of discrete-time optimization algorithms has garnered much interest since the seminal work in [1]. Continuous-time dynamics have proven useful across various engineering disciplines, including notable contributions from Hopfield and Tank in dynamical neuroscience [13], Kennedy and Chua in analog circuit design [15], and Brockett in systems and control theory [5]. There are several compelling reasons to investigate continuous-time counterparts of discrete optimization methods. First, continuous-time dynamics can offer insights into the design and convergence properties of their discrete-time equivalents. Second, differential equations that solve optimization problems in continuous time can be implemented on analog circuits [20, 15], which can lead to faster computational solutions. Finally, since biological systems inherently operate in continuous time, designing algorithms in this context opens the door to biologically plausible models that could be implemented via neural circuits. This approach may offer a deeper understanding of brain function [13, 6].
Related Work: The analysis of algorithms through the lens of systems theory is becoming popular [9]. Proximal gradient algorithms have been extensively studied in the discrete-time optimization setting [16]. A key result from this analysis is the identification of the Polyak-Łojasiewicz (PL) condition, initially introduced for gradient descent in [17] and later extended to proximal methods in [14]. When the cost function satisfies this condition in discrete time, proximal gradient algorithms achive linear convergence. The authors of [14] also discuss several existing conditions in literature and their relationships with the PL condition. Notably, they propose proximal versions of the PL and Kurdyka Łojasiewicz (KL) conditions and show their equivalence.
The study of optimization problems as continuous-time dynamical systems has a long history [1, 7, 18, 8]. The continuous-time proximal gradient dynamics were first proposed in [12] where it was shown that the PL condition implies the exponential convergence of the forward-backward envelope, even in the absence of strong convexity.
Contributions: In this work, we analyze the convergence properties and applications of proximal gradient dynamics for optimization problems with cost functions that combine a nonsmooth convex term and a potentially nonconvex smooth term. Our first contribution demonstrates that the cost function decreases monotonically along the trajectories of the proximal gradient dynamics, even in a general nonconvex setting. We then identify a natural condition that ensures exponential convergence in continuous time; we show that our condition implies the proximal KL condition. Moreover, we examine the PL condition for proximal gradient dynamics and prove that it guarantees exponential convergence of the cost function. Additionally, we derive Lipschitz-based bounds for tracking the optimal function value in parameter-varying problems under the PL condition. Finally, we explore applications of these results in sparse reconstruction, quadratic optimization, and the analysis of some matrix based problems.
The structure of the letter is as follows: Section II covers preliminaries and formulates the optimization problem. Section III presents our main result on the convergence of the flow dynamics. The proximal PL condition is discussed in Section IV, followed by applications of our results in Section V.
II Problem formulation and Preliminaries
[…]
Next, we recall some definitions.
Definition 1 (Dini Derivative):
For $\varphi\colon(a,b)\rightarrow\real$, we define the upper right Dini derivative at $t\in(a,b)$ as
$D^{+}\varphi(t):=\limsup_{h\to 0^{+}}\frac{\varphi(t+h)-\varphi(t)}{h}.$ (1)
Definition 2 (L-smoothness):
[…]
$\|\nabla f(x)-\nabla f(x)\|\leq L\|x-y\|.$ (2)
In this work, we consider the optimization problem,
$\min_{x}f(x)+g(x)=F(x),$ (3)
where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is continuously differentiable and $g:\mathbb{R}^{n}\rightarrow\overline{\real}$ is a convex closed proper (CCP) function which is potentially nondifferentiable. We begin by introducing the proximal operator and the proximal gradient dynamics.
II-A Proximal operator and the proximal gradient dynamics
The proximal operator of a closed, convex, and proper function $g\colon{}^{n}\rightarrow\overline{\real}$ is defined as
$\mathrm{prox}_{\alpha g}(x)=\underset{u\in{}^{n}}{\textup{arg\,min}}\bigl{(}\alpha g(u)+\frac{1}{2}\|u-x\|_{2}^{2}\bigr{)}.$ (4)
The proximal gradient dynamics proposed in [12], are given by,
$\dot{x}=-x+\mathrm{prox}_{\alpha g}\bigl{(}x-\alpha\nabla f(x)\bigr{)}=F_{\mathrm{prox}}(x).$ (5)
The authors of [12] prove that under assumptions on the strong convexity and smoothness of $f$, these dynamics are globally exponentially converge to the solution of optimization problem (3), as a consequence of the nonexpansive nature of the proximal operator.
II-B Polyak-Łojaciewicz and related conditions
The PL condition [17] was initially proposed to show linear convergence in discrete time setups for gradient descent.
Definition 3 (PL condition):
A differentiable function $f\colon{}^{n}\rightarrow\real$ which has a minimum at $x^{\star}\in\mathbb{R}^{n}$ is said to satisfy the PL condition if there exists a $\mu>0$ such that,
$\frac{1}{2}\|\nabla f(x)\|^{2}\geq\mu\bigl{(}f(x)-f(x^{\star})\bigr{)}.$ (6)
The PL condition is weaker than the usual assumption of convexity. A detailed discussion of the different conditions for convergence of gradient descent in discrete time is discussed in [14]. Further [14], also proposes an extension of the PL condition for the proximal case, under the assumption of $L-$smoothness.
Definition 4 (Proximal PL Condition [14]):
$f\colon{}^{n}\rightarrow\real$ and $g\colon{}^{n}\rightarrow\overline{\real}$ are said to satisfy the proximal PL condition if there exists a ${\mu>0}$ and ${\alpha>0}$, for all $x$, such that,
$\frac{1}{2}\mathcal{D}_{g}(x,\alpha)\geq\mu(F(x)-F(x^{\star})),$ (7)
where $F(x^{\star})$ is the minimum value of $F$, and
$\mathcal{D}_{g}(x,\alpha)=$
$\frac{2}{\alpha}\max_{y\in{}^{n}}\left\{\langle\nabla f(x),x-y\rangle-\frac{1}{2\alpha}\|x-y\|_{2}^{2}+g(x)-g(y)\right\}.$
Remark 1:
Note that the original definition in [14] enforces $\alpha=1/L$, where $L$ is the smoothness parameter of $f$. We relax this constraint in order to achieve a more general bound.
Definition 5 (Proximal KL condition [14]):
The KL condition with exponent $\frac{1}{2}$ holds if there exists $\hat{\mu}>0$ such that for all $x$,
$\min_{s\in\partial F(x)}\|s\|_{2}^{2}\geq 2\hat{\mu}(F(x)-F(x^{\star})),$ (8)
where $F(x)=f(x)+g(x)$ and the $\partial F(x)$ denotes the set $\{\nabla f(x)+v\,|\,v\in\partial g(x)\}$.
Lemma 2 (Equivalence of proximal PL and KL Conditions [14]):
Let $L$ be the smoothness parameter of $f$. The proximal PL condition (7) with $\alpha=1/L$, and the proximal KL Condition (8) with exponent $1/2$ are equivalent to each other, i.e., if a function $F$ satisfies either condition for some parameter value, there is guaranteed to exist a valid parameter value for which the other condition also holds.
III Global convergence of the proximal gradient dynamics
We begin by introducing a technical lemma that provides a closed-form expression for a subgradient of a CCP function evaluated at its proximal operator.
Lemma 3 (Proximal optimality condition):
Let $g\colon{}^{n}\rightarrow\overline{\real}$ be CCP. Then for $x\in{}^{n},\alpha>0$, define $z:=\mathrm{prox}_{\alpha g}(v)$. Then
$\alpha^{-1}\bigl{(}v-z\bigr{)}\in\partial g(z).$ (9)
  •       Proof:
    By definition of $z$ and $\mathrm{prox}$,
    $z$ $=\mathrm{prox}_{\alpha g}(v)=\underset{u\in{}^{n}}{\textup{arg\,min}}\;g(u)+\frac{1}{2\alpha}\|u-v\|_{2}^{2}$
    $\iff 0\in\partial g(z)+\alpha^{-1}(z-v).$
    The result follows. ∎
Next, we present our first main result which shows that under the proximal gradient dynamics, the cost function decreases along trajectories of the system. First, we present a simplified version of the result that assumes $g$ is finite-valued, in order to build intuition and provide some additional results in this special case. Then we extend the theorem to the case of a general, possibly non-finite-valued, $g$.
Theorem 4 (Nonincreasing cost function under proximal gradient dynamics):
For the optimization problem (3), let the following assumptions hold true:
  1. (A1)
    $f$ is continuously differentiable.
  2. (A2)
    $g$ is convex, closed, and proper (CCP), and takes finite values for all finite $x$.
  3. (A3)
    A global minimizer $x^{\star}$ exists, and $f(x^{\star})+g(x^{\star})=F^{*}$.
Then, for the proximal gradient dynamics (5):
  1. (i)
    the function $V(x)=f(x)+g(x)-F^{*}$ is nonincreasing along solutions and $D^{+}V(x)\leq-\frac{1}{\alpha}\|\operatorname{F_{\mathrm{prox}}}(x)\|_{2}^{2}$;
  2. (ii)
    each trajectory converges to the set of stationary points
[…]
  3. (iii)
    if additionally, there exists $\mu>0$ such that for every $x$,
    $\frac{1}{2}\|\operatorname{F_{\mathrm{prox}}}(x)\|_{2}^{2}\geq\mu\alpha^{2}(F(x)-F^{*}),$ (10)
    then each trajectory $x(\cdot)$ satisfies
    $F(x(t))-F^{*}\leq\operatorname{e}^{-2\mu\alpha t}(F(x(0))-F^{*}),$
    in other words, $F(x(\cdot))$ globally exponentially converges to $F^{*}$ with rate $02\mu\alpha>$ 0.
  •       Proof:
    For the proximal gradient dynamics (5), let $\phi_{h}\colon{}^{n}\rightarrow{}^{n}$ denote the $h$-time flow map of the dynamics. Under the assumption of continuously differentiable $f$, $\operatorname{F_{\mathrm{prox}}}$ is locally Lipschitz, and thus the flow map $\phi_{h}$ exists for sufficiently small $h\in\real$ forward in time. We consider the function $V(x)=F(x)-F^{*}$, and consider its Dini derivative with respect to time, denoting $x(t)$ by $x$ for brevity, setting ${z=\mathrm{prox}_{\alpha g}(x-\alpha\nabla f(x))}$,
    $D^{+}V(x)$ $=\limsup_{h\to 0^{+}}\frac{V(\phi_{h}(x))-V(x)}{h}$
    $\leq\nabla f(x)^{\top}(z-x)+\limsup_{h\to 0^{+}}\frac{g(\phi_{h}(x))-g(x)}{h}$
    $\leq\nabla f(x)^{\top}(z-x)+\limsup_{h\to 0^{+}}\frac{s_{x}^{\top}(x-\phi_{h}(x))}{h}$
    $=\nabla f(x)^{\top}(z-x)+s_{x}^{\top}(z-x),$ (11)
    for $s_{x}\in\partial g(x)$. By convexity of $g(x)$, we have,
    $D^{+}V(x(t))\leq\nabla f(x)^{\top}(z-x)+g(z)-g(x).$
    From convexity of $g$, we also have, $g(z)\leq g(x)-s^{\top}(x-z)$, for all $s\in\partial g(z)$. Using this, we obtain,
    $D^{+}V(x(t))$ $\leq\nabla f(x)^{\top}(z-x)+g(x)-s^{\top}(x-z)-g(x)$
    $=\nabla f(x)^{\top}(z-x)-s^{\top}(x-z).$
    Upon an application of Lemma 3 with $v=x-\alpha\nabla f(x)$, we may choose ${s=-\alpha^{-1}(z-(x-\alpha\nabla f(x)))}$. Therefore,
    $D^{+}V(x(t))$
    $\leq\nabla f(x)^{\top}(z-x)+\alpha^{-1}(z-(x-\alpha\nabla f(x)))^{\top}(x-z)$
    $=-\frac{1}{\alpha}\|x-z\|_{2}^{2}=-\frac{1}{\alpha}\|\operatorname{F_{\mathrm{prox}}}(x)\|_{2}^{2}.$ (12)
    This proves statement (i). Applying a generalization of LaSalle’s invariance principle, extended to Dini derivatives, the system converges to the set where $D^{+}V(x)=0$, which may be characterized by the set $\{x\,|\,x=z\}$. Upon an application of Lemma 3 with $v=x-\alpha\nabla f(x)$, we can see that $\{x\,|\,x=z\}=\Omega$.
    The final statement follows from substituting (10) in (12), and using Grönwall’s lemma. ∎
Remark 5:
Our choice of scaling $\mu$ in Equation (10) is motivated by an observation that is a natural extension to the existing PL condition (6) for gradient descent. Let us consider the gradient flow given by $\dot{x}=-\alpha\nabla f(x)=\operatorname{F_{gd}}(x)$. The PL condition for gradient descent presented in Definition 3 can be rewritten as,
$\frac{1}{2}\|\operatorname{F_{gd}}(x)\|^{2}\geq\mu\alpha^{2}(f(x)-f(x^{\star})).$
It is easy to see that under the gradient flow, the PL condition forces an exponential convergence rate of $2\mu\alpha$. We note that both the condition in terms of the flow and the rate are exactly equivalent to what we obtain in the proximal case.
Remark 6:
Unlike their discrete time counterparts, the proximal gradient dynamics are guaranteed to converge for all $\alpha>0$. In the discrete case, $\alpha$ is forced to be in $(0,1/L]$. We also do not assume that $f$ is $L-$smooth in our proof.
Lemma 7:
Under the following assumptions,
  1. (i)
    $f$ is continuously differentiable.
  2. (ii)
    $g$ is CCP, takes finite values for all finite $x$.
The condition (10) implies the KL condition (8), and by extension also implies the PL condition (7).
  •       Proof:
    For every $x$, consider, ${s^{\top}(-x+z)}$, where ${z=\mathrm{prox}_{\alpha g}(x-\alpha\nabla f(x))}$, and $s_{x}\in\partial F(x)$. Following the arguments in the proof of Theorem 4 from (11) to (12), we have
    $s_{x}^{\top}(-x+z)\leq-\frac{1}{\alpha}\|x-z\|_{2}^{2}$
    $\quad\implies\|s_{x}\|_{2}\|x-z\|_{2}\geq\frac{1}{\alpha}\|x-z\|_{2}^{2},$
    where the inequality follows through Cauchy Schwarz. Note that $x=z$ only occurs at the optimal values of $x$, i.e, where $F(x)=F^{*}$. In the case where $F(x)=F^{*}$, $\partial F(x)=0$ through an optimality argument. Therefore the KL condition with exponent half is true. In the case where $x\neq z$, we have for every valid $s_{x}\in\partial F(x)$,
    $\|s_{x}\|_{2}\geq\min_{s\in\partial F(x)}\|s\|_{2}\geq\frac{1}{\alpha}\|x-z\|_{2}.$
    The KL condition follows upon squaring both sides and substituting (10), and the PL condition follows due to the equivalence in Lemma 2. ∎
In Theorem 4, we show that costs along the trajectories of the proximal gradient dynamics are decreasing, in cases where $g$ is finite-valued. In many practical examples, for instance, in the case where $g$ is an indicator function, this assumption is not met. However, in Theorem 9, we will show that the cost remains non-increasing even for non-finite-valued $g$
For the proximal gradient dynamics (5), recall that $\phi_{h}\colon{}^{n}\rightarrow{}^{n}$ denotes the $h$-time flow map of the dynamics. We consider the evolution of $F(x)=f(x)+g(x)$ along trajectories of the system by studying the evolution of $t\mapsto F(\phi_{t}(x))$ for all $x$. First, we introduce an auxillary result.
Lemma 8 (Forward-invariance of $\operatorname{Dom}g$):
The domain of $F$ is forward-invariant under the proximal gradient dynamics (5). In other words, for all ${t\geq 0}$, ${x\in\operatorname{Dom}F}$ implies ${\phi_{t}(x)\in\operatorname{Dom}F}$.
  •       Proof:
    The proof is a consequence of Nagumo’s theorem since $\mathrm{prox}_{\alpha g}(x-\alpha\nabla f(x))\in\operatorname{Dom}g$ for all $x\in{}^{n}$, and $\operatorname{Dom}g=\operatorname{Dom}F$. ∎
Now, we are ready to discuss the convergence of $F(\phi_{t}(x))$ in the case where $g$ takes infinite values.
Theorem 9 (Nonincreasing cost for proximal gradient dynamics with infinite g):
Let $f\colon{}^{n}\rightarrow\real$ be continuously differentiable and $g\colon{}^{n}\rightarrow\overline{\real}$ be CCP. Let $x\in{}^{n}$ and let $\phi_{t}(x)$ be the $t$-time flow map of the proximal gradient dynamics (5). Then the map $t\mapsto F(\phi_{t}(x))$ is nonincreasing. Further, it is either always equal to +$\infty$+ ∞, is always finite, or is infinite for a finite amount of time and then becomes finite.
  •       Proof:
For a lower semicontinuous function, we can establish that it is nonincreasing provided that its upper right Dini derivative is nonpositive [10, Theorem 1.13]. We note that $F(\phi_{t}(x))$ is lower semicontinuous, as $g$ is CCP and hence lower semicontinuous, which implies $F$ is as well, which in turn implies that the composition $t\mapsto F(\phi_{t}(x))$ is lower semicontinuous. Now, we analyse each case separately. If $F(\phi_{t}(x))=+\infty$ for all $t\geq 0$, there is nothing to prove. Suppose alternatively that it is always finite. Then we know that the map $F(\phi_{t}(x))$ is lower semicontinuous and we can compute its Dini derivative, and show that it is nonpositive. Upon noting that the constant $F^{*}$ does not affect the Dini derivative $D^{+}F(\phi_{t}(x))$, we may use the same arguments obtained in Theorem 4, to obtain,
$D^{+}F(\phi_{t}(x))$ $\leq-\frac{1}{\alpha}\|\phi_{t}(x)-z\|_{2}^{2}\leq 0.$
In other words, when $F(\phi_{t}(x))$ is finite, it is nonincreasing. In the case that it is infinite and then becomes finite after some time horizon, the same argument as above applies on the interval of time when $F(\phi_{t}(x))$ is finite. Due to the forward-invariance of $\operatorname{Dom}g$, $F(\phi_{t}(x))$ is either always equal to +$\infty$+ ∞, is always finite, or is infinite for a finite amount of time and then becomes finite. In other words, once $F(\phi_{t}(x))$ is finite, it can never become +$\infty$+ ∞ again. ∎
Remark 10:
Due to the arguments made in this theorem, in the special case where $g$ is the indicator of a convex set, if the initial state is within the feasible set, then trajectories of proximal gradient dynamics always remain within the feasible set. On the other hand, if $g$ is the indicator function of the origin, then the proximal gradient dynamics are ${\dot{x}=-x}$ and for any $x\neq 0$, $F(\phi_{t}(x))$ is infinite for all $t\geq 0$ even though $\phi_{t}(x)\to 0$ as $t\to\infty$.
IV Polyak-Łojaciewicz conditions for proximal gradient dynamics
Next, we discuss sufficient conditions for exponential convergence of the cost function of (3) under proximal gradient dynamics (5).
Theorem 11 (Exponential convergence under the PL condition):
For the optimization problem (3) under the proximal gradient dynamics (5), such that $f(x)+g(x)$ satisfies the proximal PL condition (7) with constants $\mu>0,\alpha>0$ and $g$ is CCP, we have guaranteed global exponential convergence to the minimizer $F^{*}$ with some rate $c={\mu\alpha}>0$, that is,
$F(x(t))-F^{*}\leq\operatorname{e}^{-ct}(F(x(0))-F^{*}).$
  •       Proof:
    Note that if $V(x(0))$ is infinity, there is nothing to show. If $V(x(0))$ is finite, it stays finite for all future timesteps. Similar to the proof of Theorem 4, consider the Dini derivative of ${V(x(t))=f(x(t))+g(x(t))-F^{*}}$. Using the arguments presented in Theorem 4 upto Equation (11),
    $D^{+}V(x)\leq\nabla f(x)^{\top}(z-x)+s_{x}^{\top}(-x+z).$
    where $z=\mathrm{prox}_{\alpha g}(x-\alpha\nabla f(x))$, and $s_{x}\in\partial g(x)$. From the convexity of $g$, we have,
    $D^{+}V(x)\leq\nabla f(x)^{\top}(z-x)+g(z)-g(x).$
    Adding and subtracting $\frac{1}{2\alpha}\|x-z\|_{2}^{2}$,
    $D^{+}V(x)$ $\leq\nabla f(x)^{\top}(z-x)+g(z)-g(x)$
    $\quad+\frac{1}{2\alpha}\|x-z\|_{2}^{2}-\frac{1}{2\alpha}\|x-z\|_{2}^{2}.$ (13)
    Note that
[…]
    $\quad+\langle\nabla f(x),u-x\rangle\bigr{\}}$
[…]
    Therefore,
    $\nabla f(x)^{\top}(z-x)+g(z)-g(x)+\frac{1}{2\alpha}\|x-z\|_{2}^{2}$
    $\quad=\min_{y}\left(\nabla f(x)^{\top}(y-x)+g(y)-g(x)+\frac{1}{2\alpha}\|x-y\|_{2}^{2}\right)$
    $\quad=-\frac{\alpha}{2}\mathcal{D}_{g}(x,\alpha)$
    Using this equality in (Proof:),
    $D^{+}V(x)$ $\leq-\frac{\alpha}{2}\mathcal{D}_{g}(x,\alpha)-\frac{1}{2\alpha}\|x-z\|_{2}^{2}$ (14)
    Using the proximal PL condition (7),
    $D^{+}V(x)$ $\leq-\alpha\mu(F(x)-F(x^{\star}))-\frac{1}{2\alpha}\|x-z\|_{2}^{2}$ (15)
    Now, the second term may be bounded above by 00. Using Grönwall’s lemma we get our required result. ∎
Remark 12:
The PL condition in continuous time is only of value when $V(x(0))$ is finite. In discrete time, the proximal gradient algorithm forces a finite value for $V(\cdot)$ in one timestep. In general, in continuous time, one cannot characterize the time taken for $V(x(\cdot))$ to become finite.
Remark 13:
In [12, Theorem 3], the forward-backward envelope is shown to be globally exponentially stable under the PL conditions. The forward-backward envelope is an upper bound to $f(z)+g(z)$, under the assumption of $L-$smoothness of $f$, where $z=\mathrm{prox}_{\alpha g}(x-\alpha\nabla f(x))$. In contrast, here we are interested in the value of $f(x)+g(x)$. Identifying an upper bound on $f(z)+g(z)$ is particularly useful when $g(\cdot)$ takes nonfinite values but also does not reflect whether the current state is feasible.
IV-A Extensions to time-varying optimization
In recent literature [8], variations of the classic problem (3) with time-varying parameters have also been considered. In this case, the goal is to define dynamics that track the minimizer of the optimization problem,
$\min_{x}f(x,\theta(t))+g(x,\theta(t))=F(x,\theta(t)).$ (16)
Under Lipschitz constraints on $\theta$, one may define bounds for the difference of the current cost and minimum cost, when the PL condition is true.
Theorem 14 (Cost tracking under the PL condition):
Consider the time-varying optimization problem (16). Assume $f(x,\theta(t))$ and $g(x,\theta(t))$ satisfy the proximal PL condition (7) with constant $\mu$ at each time $t$, $g$ is CCP, and $F(x,\theta)$ is differentiable with respect to, and Lipschitz continuous in $\theta$, with parameter $\ell_{\theta}$. Given an initial condition $x(0)$ and a parameter trajectory $\theta(t)$, define ${V(t)=F(x(t),\theta(t))-F(x^{\star}(t),\theta(t))}$. Then, under the proximal gradient dynamics (5)
  1. (i)
    $D^{+}V(t)\leq-\mu\alpha V(t)+2\ell_{\theta}\|\dot{\theta}(t)\|$, and
  2. (ii)
    the Grönwall inequality for Dini derivatives implies
    $V(t)\leq\operatorname{e}^{-\mu\alpha t}V(0)+\frac{2\ell_{\theta}}{\mu\alpha}\int_{0}^{t}\operatorname{e}^{-\mu\alpha(t-\tau)}\|\dot{\theta}(\tau)\|d\tau.$
  •       Proof:
    Through arguments similar to the proofs of Theorems 4 and for any $s_{1}\in\partial_{x}F(x,\theta),s_{2}\in\partial_{x}F(x^{\star},\theta)$,
    $D^{+}V(x,t)$ $\leq s_{1}^{\top}\dot{x}+(\nabla_{\theta}F(x,\theta)-\nabla_{\theta}F(x^{\star},\theta))^{\top}\dot{\theta}$
    $\quad-s_{2}^{\top}\dot{x}^{\star}(t).$ (17)
    Following the arguments of Theorem 11, the first term is bounded above by $-\alpha\mu(V(x,t))$. Since $F$ is Lipschitz in $\theta$, $\|\nabla_{\theta}F(x)\|\leq\ell_{\theta}$. Through an application of Cauchy Schwarz and triangle inequality, the second term may be bounded by $2\ell_{\theta}\|\dot{\theta}\|$ The final term is 00 due to the optimality of $x^{\star}$. In summary, we have,
    $D^{+}V(x,t)\leq-\mu\alpha V(x,t)+2\ell_{\theta}\|\dot{\theta}\|.$
    Using Grönwall’s lemma, we obtain the second result. ∎
V Applications and Examples
In this section, we discuss a few applications of the proximal gradient dynamics.
V-A Least Squares with $\ell_{1}$ regularization
Through the proximal gradient dynamics, one may solve the $\ell_{1}$ regularized least squares problem, otherwise known as the Least Absolute Shrinkage and Selection Operator (LASSO). The problem is given by,
$\min_{x\in\mathbb{R}^{n}}\frac{1}{2}\|Ax-u\|_{2}^{2}+\lambda\|x\|_{1},$ (18)
where $A\in\mathbb{R}^{m\times n},\;u\in\mathbb{R}^{m}$ The corresponding proximal gradient dynamics, referred to as the “firing rate competitive network” [6] are given by,
$\dot{x}=-x+\operatorname{soft}_{\alpha\lambda}\bigl{(}{(I_{n}-\alpha A^{\top}A)x+\alpha A^{\top}u}\bigr{)}.$ (19)
Here, $\operatorname{soft}_{\alpha\lambda}({\cdot})$ represents the soft thresholding operator and is the proximal operator of the $\ell_{1}$ norm:
$\operatorname{soft}_{\alpha\lambda}({x_{i}})=\begin{cases}x_{i}-\alpha\lambda\operatorname{sign}({x_{i}})&\text{if}\quad|x_{i}|>\alpha\lambda,\\ 0&\text{if}\quad|x_{i}|\leq\alpha\lambda\end{cases}$
The paper [6] also includes a bound on the convergence to the optimal value of $x$ under the restricted isometry property.
Corollary 15:
Under the proximal gradient dynamics (19) for the LASSO problem (18), the cost function decays to the optimal value exponentially.
  •       Proof:
    The LASSO problem is known to satisfy the PL condition [14]. The result follows from Theorems 11 and 14. ∎
V-B Quadratic optimization with polytopic constraints
Next, we consider another classic problem that is of interest in several domains. Consider the optimization problem,
$\min_{x\in\mathbb{R}^{n}}\frac{1}{2}x^{\top}Qx-u^{\top}x\quad\text{s.t.}\quad Ax\leq b,\;Dx=c.$ (20)
where $Q\in\mathbb{R}^{n\times n}$ is positive definite, and other matrices are appropriately sized. Specific versions of this problem have been implemented both as neural circuits [4] and using analog circuits [20]. Further, it also represents the Fenchel dual to the cost function of support vector machines (SVMs) [14]. The corresponding proximal gradient dynamics are given by,
$\dot{x}=-x+\mathrm{proj}_{C}(x-\alpha Qx+\alpha u),$ (21)
where $C=\{x\;|\;Ax\leq b,\;Dx=c\}$.
Corollary 16:
Under the proximal gradient dynamics (21) for the constrained quadratic optimization problem (20), if the initial point is within the feasible set, and $\alpha=1/\lambda_{\min}(Q)$, the cost function decays to the optimal value exponentially with rate 1.
  •       Proof:
    Since we begin in the feasible set, the initial cost is finite. Therefore, by Theorem 9, the cost function converges. Since $f(x)$ is strongly convex, and the constraints set is convex, the proximal PL condition is true [14], with $\alpha=1/\lambda_{\min}(Q)$, and $\mu=\lambda_{\min}(Q)$. By applying Theorem 11, the rate of convergence for the cost function is $1$. ∎
V-C Matrix based problems
Finally, we consider matrix optimization problems. In this setting, the analogue of sparse vectors are low rank matrices. However, similar to the $\ell_{0}$ norm, imposing a low rank constraint leads to nonconvexity. A known convex relaxation for the $\ell_{0}$ norm is the trace norm. The proximal operator for $g(X)=\|X\|_{*}$ is
$\mathrm{prox}_{\alpha g}(X)=\operatorname{S}_{\alpha}({X})=U\Sigma_{\alpha}V^{\top},$ (22)
where $X\in\mathbb{R}^{m\times n},X=U\Sigma V^{\top}$ is an SVD and $\Sigma_{\alpha\lambda}$ is diagonal with $(\Sigma_{\alpha\lambda})_{ii}=\operatorname{soft}_{\alpha\lambda}({\Sigma_{ii}})$ [2]. A general class of problems including matrix recovery and matrix completion problems can be written as
$\min_{X\in\mathbb{R}^{m\times n}}\frac{1}{2}\|y-\mathcal{A}[X]\|_{2}^{2}+\lambda\|X\|_{*}$ (23)
where ${\mathcal{A}[X]=[\operatorname{tr}(A_{1}^{\top}X),\operatorname{tr}(A_{2}^{\top}X),\cdots\operatorname{tr}(A_{k}^{\top}X)]^{\top}}$, with each $A_{i}\in\mathbb{R}^{m\times n}$ and ${y\in\mathbb{R}^{k}}$. The proximal gradient dynamics for this problem are
$\dot{X}=-X+\operatorname{S}_{\alpha\lambda}\bigl{(}{X-\alpha\sum_{i=1}^{k}(\operatorname{tr}(A_{i}^{\top}X)-y_{i})A_{i}}\bigr{)},$ (24)
A second popular matrix based optimization problem is the matrix factorization problem. This problem and its generalizations appear in overparameterized linear neural networks with mean square error loss [21]. In general, this problem is nonconvex. To promote a low-rank solution, it is customary to add terms proportional to the nuclear norm and write the problem as:
$\min_{W_{1},W_{2}}\frac{1}{2}\|Y-XW_{1}W_{2}\|_{F}^{2}+\lambda\|W_{1}\|_{*}+\lambda\|W_{2}\|_{*},$ (25)
where $Y\in\mathbb{R}^{N\times m},X\in\mathbb{R}^{N\times n},W_{1}\in\mathbb{R}^{n\times h}$ and $W_{2}\in\mathbb{R}^{h\times m}$. The proximal gradient dynamics are
$\dot{W}_{1}=-W_{1}+\operatorname{S}_{\alpha\lambda}\bigl{(}{X-\alpha X^{\top}(Y-XW_{1}W_{2})W_{2}^{\top}}\bigr{)},$ (26a)
$\dot{W}_{2}=-W_{2}+\operatorname{S}_{\alpha\lambda}\bigl{(}{X-\alpha W_{1}^{\top}X^{\top}(Y-XW_{1}W_{2})}\bigr{)}$ (26b)
Both (24) and (26) are guaranteed to converge to a minimizer or a saddle point via Theorem 4. It is an open question whether the costs in (23) and (25) are exponentially decreasing along trajectories of (24) and (26), respectively.
VI Discussion
In this letter, we discuss some convergence properties of the proximal gradient dynamics. We prove in Theorem 4 that the cost function of the associated optimization problem is nonincreasing for finite-valued regularizers. For non-finite-valued regularizers, we are able to show that the cost is nonincreasing in Theorem 9. Next, we extend the PL conditions to continuous-time proximal gradient dynamics, explicitly showing that the cost function decays exponentially under the PL condition. We use this result to consider the case of parametric optimization problems with time-varying parameters, and provide a cost function tracking bound.
An interesting direction of future work is the application of proximal gradient dynamics to design continuous-time solvers, such as analog circuits and biologically plausible neural networks. Here, we discuss constrained optimization, LASSO problem and matrix recovery and factorization. However, these dynamics may also allow a generalized nonsmooth analysis of systems based on energy landscapes. A system that is currently studied under this framework is the human brain [11]. While previous work [19] suggests that the algorithms we design for modeling a brain must be gradient-based, due to these dynamics, one could design biologically plausible algorithms for a nonsmooth energy landscape.